OpenStudy (anonymous):

In how many ways can three different numbers be selected from the numbers 1 -12, so that their sum can be exactly divided by 3? Please give a good way of doing it. Thanks

OpenStudy (joannablackwelder):

Every third number would be divisible by 3. i.e. 1+1+1=3, 1+1+2=4, 1+1+3=5 etc.

OpenStudy (joannablackwelder):

Do you know how to find the total number of numbers?

OpenStudy (anonymous):

That's what I want to know.

OpenStudy (joannablackwelder):

Each number has 12 options and there are three numbers. So 12^3.

OpenStudy (wolf1728):

The number of combinations in which 3 (k) things can be selected from 12 (n) things equals n! / k! • (n - k)! = 12! / 3! • (9!) = 12 • 11 • 10 • 9! / 6 • 9! = 2 • 11 • 10 = 220 ways in which 3 numbers can be selected from 12 These would go from (1, 2, 3) all the way to (10, 11, 12) If the sum of these 3 numbers is divisible by 3, 6, 9, 12, 15, etc then the sum is divisible by 3. So, this narrows down the choices even further to 220/3 or 73 groups of numbers. And heck, if you want to know what those 73 groups of numbers are, just look here: 1 2 3 1 2 6 1 2 9 1 2 12 1 3 5 1 3 8 1 3 11 1 4 7 1 4 10 1 5 6 1 5 9 1 5 12 1 6 8 1 6 11 1 7 10 1 8 9 1 8 12 1 9 11 1 11 12 2 3 4 2 3 7 2 3 10 2 4 9 2 4 12 2 5 8 2 6 7 2 6 10 2 7 9 2 7 12 2 8 11 2 9 10 2 10 12 3 4 5 3 4 8 3 4 11 3 5 7 3 5 10 3 6 9 3 6 12 3 7 8 3 7 11 3 8 10 3 9 12 3 10 11 4 5 6 4 5 9 4 5 12 4 6 8 4 6 11 4 7 10 4 8 9 4 8 12 4 9 11 4 11 12 5 6 10 5 7 9 5 7 12 5 8 11 5 9 10 5 10 12 6 7 8 6 7 11 6 8 10 6 9 12 6 10 11 7 8 9 7 8 12 7 9 11 7 11 12 8 9 10 8 10 12 9 10 11 10 11 12

OpenStudy (anonymous):

Thanks very much

Parth (parthkohli):

Do you know a little bit of modular arithmetic?